OpenStudy (anonymous):

Find an exact value. sine (-11pi/12)

OpenStudy (anonymous):

is it\[(-\sqrt{6}-\sqrt{2})\div4\]

OpenStudy (jdoe0001):

hmmm no exactly...

OpenStudy (jdoe0001):

it's just a angle sum

OpenStudy (jdoe0001):

let's take a peek at the rational only -11/12 how can we split it in say, 2 fractions and have the result be -11/12

OpenStudy (jdoe0001):

anyhow... so I used.. 9/12 - 20/12

OpenStudy (jdoe0001):

pretty simple thus far

OpenStudy (debbieg):

Your answer is almost correct, I think you just have a sign error in there somewhere.

OpenStudy (anonymous):

\[\sqrt{2}-\sqrt{6}/4\]

OpenStudy (anonymous):

is that it?

OpenStudy (jdoe0001):

dunno

OpenStudy (jdoe0001):

what you need is => http://hyperphysics.phy-astr.gsu.edu/hbase/imgmth/trid1.gif <-- the sum identities

OpenStudy (jdoe0001):

so if I were to used 9/12 - 20/12 = -11/12 well then that leaves me with \(\bf sin\left(-\cfrac{11\pi}{12}\right)\\\quad \\ \cfrac{9\pi}{12} - \cfrac{20\pi}{12} \implies \cfrac{3\pi}{4} - \cfrac{5\pi}{3}\quad \\\quad \\ sin\left(-\cfrac{11\pi}{12}\right) \implies sin\left(\cfrac{3\pi}{4} - \cfrac{5\pi}{3}\right)\)

OpenStudy (jdoe0001):

and those are the angles you can use to add up, using the trig sum identities

OpenStudy (jdoe0001):

do you see the sum identity you need in the picture?

OpenStudy (anonymous):

what the answer though

OpenStudy (jdoe0001):

the answer is, whatever the sum gives you you get the values for the angles from the Unit Circle

OpenStudy (jdoe0001):

keep in mind the idea behind exercises is not to transcribe answers is to do them so one can understand it'd be of no use if a teacher hands out tests and also hands out the answers alonside, so students can just transcribe them

OpenStudy (anonymous):

\[\sin \left( \frac{ -11\pi }{ 12 } \right)=-\sin \frac{ 11\pi }{ 12 }=-\sin \left( \pi-\frac{ \pi }{ 12 } \right) \[=-\sin \frac{ \pi }{ 12 }\] \[=-\sin \left( \frac{ \pi }{ 3 }-\frac{ \pi }{ 4 } \right)\] \[=-\left[ \sin \frac{ \pi }{ 3 }\cos \frac{ \pi }{ 4 }-\cos \frac{ \pi }{3}\sin \frac{ \pi }{ 4 } \right]\] \[=-\left[ \frac{ \sqrt{3} }{2 }*\frac{ 1 }{ \sqrt{2} }-\frac{ 1 }{2 }*\frac{ 1 }{ \sqrt{2} } \right]\] \[=-\frac{ \sqrt{3}-1 }{ 2\sqrt{2} }\]

OpenStudy (debbieg):

If you rationalize that^^^ den'r, notice that you get: \(\Large =-\dfrac{ \sqrt{3}-1 }{ 2\sqrt{2} }\cdot\dfrac{ \sqrt{2} }{\sqrt{2} }=-\dfrac{ \sqrt{6}-\sqrt{2} }{ 4}=\dfrac{ -\sqrt{6}+\sqrt{2} }{ 4}\) which is ALMOST what you answered up above, but for a sign error in the 2nd term in the num'r. So you almost had it, but you should trace your error to see where you went wrong with the sign, so that you won't make that same mistake again.